OpenStudy (anonymous):

dy/dx=x^2y^2 d^2y/dx^2=?

OpenStudy (anonymous):

take the derivative of x^2 y^2 using the product rule

OpenStudy (zzr0ck3r):

use implicit differentiation when you do it

OpenStudy (anonymous):

^

OpenStudy (zzr0ck3r):

you will have a dy/dx in it, then plug your original thing back in.

OpenStudy (zzr0ck3r):

\(\frac{dy}{dx}=x^2y^2\\\frac{d^2y}{dx^2}=2xy^2+2y\frac{dy}{dx}x^2=2xy^2+2y(x^2y^2)x^2=2xy^2+2x^4y^3\)